a cyclist traveled at at rate of 32 mph to visit a nearby town. The cyclist averaged 10 mph on the return trip. If the trip took 6.3 hours, find the distance to the nearby town.
Answer by checkley77(12569)

D=RT OR T=D/R
X/32+X/10=6.3
(10X+32X)/(10*32)=6.3
42X/320=6.3
42X=6.3*320
42X=2,016
X=2,016/42
X=48 MILES IS THE DISTANCE TO THE NEARBY TOWN.
PROOF:
48/32+48/10=6.3
1.5+4.8=6.3
6.3=6.3